# Rethinking the Reliability of Multi-agent System: A Perspective from Byzantine Fault Tolerance

### v1

@@ Introduction
Inspired by these findings, we propose CP-WBFT, a confidence probe-based weighted Byzantine Fault Tolerant consensus mechanism to enhance the stability of MAS with various topologies. It leverages the inherent reflective and discriminative capabilities of LLMs to enhance the reliability of agents. Specifically, we first develop confidence probes from both the prompt and decoder perspectives to assess the agent’s confidence level. We then introduce a confidence-guided Byzantine consensus protocol that enhances the reliability of LLM-based agents by using probe-based weighted information flow, assigning higher transmission weights to more credible agents. Extensive experiments demonstrate that CP-WBFT outperforms superior performance across diverse network topologies, with HCP achieving +85.71% Byzantine Fault Tolerance Improvement on complete graphs across both mathematical reasoning and safety assessment tasks, while maintaining 100% round-level accuracy. Our main contributions are summarized as follows:
@@
  • We propose a novel probes tailored for MAS, which leverage the inherent reflective and discriminative capabilities of LLMs to effectively identify problematic agents from both prompt-level and hidden-level.
@@ Methods
Motivated by the pilot experiments demonstrating LLM-based agents’ superior Byzantine fault tolerance, we develop a comprehensive confidence probes from both the hidden- and prompt-level. It leverages the LLM agent’s inherent reflective to extract fine-grained confidence signals for enhanced Byzantine fault tolerance in MAS.
@@
Topology sensitivity analysis reveals hierarchical performance patterns across protocols. As shown in Table 3, on GSM8K, HCP consistently outperforms PCP across all network configurations, with the performance gap being most pronounced in well-connected structures like complete graphs and least evident in constrained topologies such as chains. Star topologies exhibit structure-dependent behavior: PCP demonstrates substantially answer resilience when leaves are malicious compared to when the center node is compromised, highlighting the critical role of central nodes in information propagation. On XSTest, HCP achieves perfect consensus ianswern complete graphs, while PCP shows task-specific vulnerability. Unlike its modest performance on GSM8K, PCP exhibits neutral or even negative BFTI on several XSTest topologies, particularly in tree structures and star configurations with malicious centers. This divergence underscores the heightened topology sensitivity of safety assessment tasks compared to mathematical reasoning. Figure 4 illustrates CP-WBFT’s practical workflow, demonstrating how the system achieves consensus on a mathematical reasoning problem under Byzantine faults.

### v2

@@ Introduction
Inspired by these findings, we propose CP-WBFT, a confidence probe-based weighted Byzantine Fault Tolerant consensus mechanism to enhance the stability of MAS with various topologies. It leverages the inherent reflective and discriminative capabilities of LLMs to enhance the reliability of agents. Specifically, we first develop confidence probes from both the prompt and decoder perspectives to assess the agent’s confidence level. We then introduce a confidence-guided Byzantine consensus protocol that enhances the reliability of LLM-based agents by using probe-based weighted information flow, assigning higher transmission weights to more credible agents. Extensive experiments demonstrate that CP-WBFT achieves superior performance across diverse network topologies, with HCP achieving +85.71% Byzantine Fault Tolerance Improvement on complete graphs across both mathematical reasoning and safety assessment tasks, while maintaining 100% round-level accuracy. Our main contributions are summarized as follows:
@@
  • We propose a novel probe tailored for MAS, which leverage the inherent reflective and discriminative capabilities of LLMs to effectively identify problematic agents from both prompt-level and hidden-level.
@@ Methods
Motivated by the pilot experiments demonstrating LLM-based agents’ superior Byzantine fault tolerance, we develop comprehensive confidence probes from both the hidden- and prompt-level. It leverages the LLM agent’s inherent reflective capabilities to extract fine-grained confidence signals for enhanced Byzantine fault tolerance in MAS.
@@
Topology sensitivity analysis reveals hierarchical performance patterns across protocols. As shown in Table 3, on GSM8K, HCP consistently outperforms PCP across all network configurations, with the performance gap being most pronounced in well-connected structures like complete graphs and least evident in constrained topologies such as chains. Star topologies exhibit structure-dependent behavior: PCP demonstrates substantially resilience when leaves are malicious compared to when the center node is compromised, highlighting the critical role of central nodes in information propagation. On XSTest, HCP achieves perfect consensus in complete graphs, while PCP shows task-specific vulnerability. Unlike its modest performance on GSM8K, PCP exhibits neutral or even negative BFTI on several XSTest topologies, particularly in tree structures and star configurations with malicious centers. This divergence underscores the heightened topology sensitivity of safety assessment tasks compared to mathematical reasoning. Figure 4 illustrates CP-WBFT’s practical workflow, demonstrating how the system achieves consensus on a mathematical reasoning problem under Byzantine faults.